OpenStudy (anonymous):

Solve by factoring the GCF: 16x^2-4x=0 A. {0, 1/3} B. {o, 4} C. {4, -4} D. {-1, 4}

jimthompson5910 (jim_thompson5910):

16x^2-4x=0 4x(4x-1) = 0 4x = 0 or 4x - 1 = 0 x = 0 or x = 1/4